OpenStudy (anonymous):

sin θ(2 cos θ + 1) = 0

OpenStudy (anonymous):

\[2\sin \theta \cos \theta+\sin\]

OpenStudy (anonymous):

\[2\sin^2 \theta=2\sin \theta \cos \theta\]

OpenStudy (anonymous):

2sin^2 theta +sintheta=0 again quadratic

OpenStudy (anonymous):

can u solve?